# Fairness in Social Influence Maximization via Optimal Transport

### v3

@@ 3.2 Mutual Fairness in Practice
We now investigate the use of our newly defined fairness metric across a variety of real-world datasets: Add Health (AH), Antelope Valley variants 00 to $23$ (AV_$\{0\text{-}23\}$[26], APS Physics (APS[14], Deezer (DZ[19], High School Gender (HS[15], Indian Villages (IV[1], and Instagram (INS[22]. Each dataset contains a social network with a chosen demographic dividing the population into two non-overlapping groups (see Appendix C for details). We load the datasets as graphs $G(V,E)$. We then select a seedset $S$ of size $2$-$90$ (depending on the dataset) using the following heuristics: two group-agnostic seed selection strategies as our baselines, namely degree centrality (bas_d), and greedy (bas_g), proposed by Kempe et al. [12]. In addition, we implement two fair seed selection heuristics based on the equity metric, namely degree-central fair heuristic (hrt_d), and greedy fair heuristic (hrt_g), proposed by Stoica et al. [23]. To model the information spread, we use the Independent Cascade model (IC) for the diffusion of information [12] with a probability $p\in[0,1]$ for all edges. This process, being stochastic, is simulated $R$ times in a Monte Carlo sampling process to achieve $R$ final configurations (Definition 2.1) plotted together as a joint outreach distribution, in Fig. 3. Then we apply our distribution-aware notion of fairness from Section 3.1, mutual fairness. We keep $R=1,000$ throughout, but explore several values in $p,|S|$ (mentioned per experiment in the figures below) and exhaustively recorded with other hyperparameters in Appendix D. All details related to computational resources and development environment are available in Appendix G. The code for all our numerical experiments is available at https://github.com/nicolaslanzetti/fairness-sim-ot.
@@
As a first experiment, we study the joint outreach probability distribution for different datasets. We identify four qualitatively different outcomes, shown in Fig. 3 for a few of the datasets. Additional experiments with different propagation probability and seed selection strategies can be found in Appendix D. Fig. 3(a) is obtained on AH with bas_g selection strategy and $p=0.5,|S|=10$. We note how the joint outreach distribution is almost concentrated on the top right of the plane, i.e., the outcome is almost deterministic and highly fair and efficient. In turn, this trivializes both the expected value in the equity metric and the cost in the mutual fairness metric in Definition 3.1, which therefore essentially boils down to comparing the almost deterministic outreach fraction within each group. In these cases, our fairness metric does not provide additional insights. Such deterministic outcomes are typical of degree or greedy seedset outreach in dense graphs, such as AH, DZ, INS (refer to Appendix D), with extreme probability of conduction ($p\geq 0.5$ or $p\rightarrow 0$), and cross-group interconnectivity (see Table 1 in Appendix C). For moderate $p$ (e.g., $0.1$), the outreach probability distribution is concentrated along the diagonal (Fig. 3(b)). Thus, both the equity metric and our fairness measure are maximal. Nonetheless, our fairness metric provides additional insights: not only does the expected outreach within each group coincide, but also the outreach at every realization coincides (see the example in Section 3). Thus, our fairness metric provides a stronger certificate of fairness. As before, the same applies to AH, DZ, INS (see Appendix D). Intuitively, high cross-group interconnectivity in a dense graph already ensures fairness. Additionally, extreme $p$ values ensure deterministic outreach (either the information dies out at the seedset, or reaches everyone in the population). When propagation happens with moderate propagation probabilities, $p$, outreach appears as in Fig. 3(b). Fig. 3(c) represents APS for its hrt_g seedset outreach and $p=0.3,|S|=6$. Here, we observe a highly stochastic outcome, with many realizations for which almost no member of one group receives the information. Note that the phenomenon observed in Fig. 3(c) is the same as the one captured by our motivating example. We argue such an outcome should not be classified as fair, despite the expected value of the proportions being similar. Finally, Fig. 3(d) shows the AV_0 dataset with $p=0.3,|S|=4$, and bas_g selection strategy. We observe a more stochastic outreach compared to Fig. 3(b) with variance spread along, but not on the diagonal, with a small bias towards one group. Also in this case, both the equity and the mutual fairness metrics characterize this outcome as fair, but mutual fairness is more informative as it requires outcomes to be fair at each realization.
@@
As a second experiment, we investigate the difference between mutual fairness and equity (difference in the expected value of the proportions), as a function of the conduction probability $p$. We consider the IV dataset as a case study and select seeds using bas_g. We show our results in Fig. 4. Our mutual fairness metric in Definition 3.1 shows a fundamentally different trend compared to the equity metric from Definition A.3. Importantly, for $p\in(0,0.5)$, both metrics have an opposite trend: equity fairness increases to some extent whereas our metric suggests a huge fall in fairness in this region. For $p\in(0.5,0.7)$, there is a decrease in equity fairness, while our fairness evaluation remains relatively constant. We notice similar trends for both metrics only for $p\in(0.8,1.0)$. The significant difference in trend of the two metrics confirms our previous finding that mutual fairness is more informative than the equity metric and that the equity metric fails to adequately capture changes in fairness, see Section 3.1 and Section 3.2. For more experiments on other datasets, we refer to Section D.2.
@@ 3.3 Trading off Fairness and Efficiency
In our setting, we can easily introduce the tradeoff in the transportation cost (2). Specifically, we can define the transportation cost as a weighted sum of the “diagonal distance” (measuring difference in efficiency, dotted segment in Fig. 2) and the “orthogonal distance” (measures difference in fairness, solid segment in Fig. 2). Formally, for a given weight $\beta\geq 0$, we define the transportation cost
@@
The factors $1$ and and $\sqrt{2}/\max\{1,2-2\beta\}$ in (4) ensure that the metric is non-negative and in $[0,1]$. Again, the optimal transport problem can be solved in closed form, which yields
@@ 4.1 Fairness-promoting Seed-selection Algorithm
Armed with a novel fairness metric, $\beta\mathrm{-}\textsc{Fairness}$, we now design an iterative seed-selection algorithm, which we call Stochastic Seedset Selection Descent (S3D), that strategically selects seeds taking into account all communities simultaneously. The pseudo-code is summarized in Algorithm 1. For its motivation and details, refer to Appendix E, E.3. For a given initial seedset, our algorithm explores new seeds and evaluates them on the efficiency-fairness metric $\beta\mathrm{-}\textsc{Fairness}$ as in (4) for a desired value of the fairness-efficiency tradeoff parameter $\beta$ (S3D_STEP() in Appendix E), to decide if the new seedset becomes a candidate for the optimized seedset. These seeds are searched for by iteratively sampling stochastically reachable nodes, up to a fixed depth, taken as a fraction of the graph diameter, from the current seedset (SEEDSET_REACH() in Appendix E) while making sure they contribute to a non-overlapping outreach (Algorithm 1::6-8). To prevent getting stuck at some local minima of the generally non-convex objective, the procedure allows for visiting inferior seedsets on $\beta\mathrm{-}\textsc{Fairness}$ or even selecting completely random ones on rare occasions (Algorithm 1::12-18) using Metropolis Sampling [5]. Otherwise, a high $\beta\mathrm{-}\textsc{Fairness}$ encourages opting for the new seedset with high probability. Finally, we revisit all the seedset candidates collected so far and pick the one with the largest $\beta\mathrm{-}\textsc{Fairness}$ as the optimal seedset. For a sparse graph $G(V,E)$, with $E=O(V)$, choosing $|S|$ seeds, averaging over $R$ realizations to approximate outreach via Monte-Carlo sampling and exploring $k$ candidates using S3D_STEP suggests a total running time upper bound of $O(kR|S||V|)$ (see Appendix E for details about the algorithm complexity). In practice, $k\in[500,1000],R=1000$ for $S\in[2,100]$ works well for all datasets.
@@ 4.2 Real-world Data
We test our algorithm across a variety of datasets (Appendix C) against our baselines (bas_d, bas_g). We initialize the S3D algorithm with the two baseline seedsets and hence include results from two separately optimized seedsets, S3D_d, S3D_g. Our results are shown in Fig. 6. Informally, we observe that our seed-selection mechanism “moves” the probability mass of the joint outreach probability distribution towards the diagonal, which ultimately increases the fairness of the resulting configuration. At the same time, efficiency either increases as well or suffers only a small decrease, as we investigate more in detail in our next experiment. Generally speaking, datasets with high cross-group connections (AH, DZ, INS) can already benefit a lot from label-blind seed selection to get moderately fair outreach. Similarly, for datasets with low cross-group connections (APS) a label-blind strategy, in order to maximize efficiency, selects a diverse population of seeds from which all communities are reached. Therefore, label-blind algorithms work similarly to S3D. In other moderate cases (AV, HS, IV), instead, we observe significant improvements of S3D over label-blind strategies.
@@
In our final experiments, we compare several algorithms along with ours in terms of efficiency and mutual fairness across various datasets (see Appendix C for all datasets),. We consider the following algorithms: bas_d, bas_g, their fair heuristic counterparts, hrt_d, hrt_g, against our S3D_d, S3D_g, initialized via greedy and degree centrality baseline seeds, respectively. We show our results in Fig. 7. S3D achieves in almost all cases the highest fairness score ($y$-axis) and generally a slightly lower efficiency score ($x$-axis), compared to others. Thus, our seed-selection mechanism leads to fairer outcomes with only a minor decrease in efficiency.
@@
CEis high (>50%): The case where nodes interact more across groups than in their own groups was never observed. However, as long as the existence of edges does not reliably signal group information, we expect S3D to perform well based on a similar analysis.
@@ Appendix A Existing Fairness Metrics
$\bar{x}_{i}:=\frac{\mathbb{E}[|v\text{ reached }|v\in C_{i}|]}{|C_{i}|},\quad\forall i\in\{1,\dots,m\}.$
@@
$\frac{\mathbb{E}[|v\in S|v\in C_{i}|]}{|C_{i}|}=\frac{\mathbb{E}[|v\in S|v\in C_{j}|]}{|C_{j}|}\quad\forall i,j\in\{1,\dots,m\}.$
@@
Given a network with communities $C_{1},\dots,C_{m}$, a SIM algorithm that selects a seedset $S$ is said to be equitable if the algorithm propagation reaches all communities in a balanced way, i.e. $\bar{x}_{i}=\bar{x}_{j}$, $\forall i,j\in\{1,\dots,m\}$.
@@ Appendix B Extension to Multiple Groups
We start with the definition of mutual fairness. We proceed as Section 3.1 and define mutual fairness via optimal transport, which, in turn, requires defining a reference distribution and a transportation cost. The reference distribution is again the “ideal” distribution $\gamma^{\ast}=\delta_{(1,\ldots,1)}$ which encodes the case in which all members of all groups receive the information. As for the transportation cost, it suffices to generalize the transportation cost (2) to an $m$ dimensional space. Specifically, it can be defined as the distance between any given point $(x_{1},\ldots,x_{m})\in[0,1]^{m}$ in the hypercube and the diagonal line. For this, let
@@
$\textsf{Fairness}(\gamma)$ $=1-\alpha\mathbb{E}_{(x_{1},\ldots,x_{m})\sim\gamma}[\norm{z(x_{1},\ldots,x_{m})-(x_{1},\ldots,x_{m})}]$
@@
where the constant $\alpha>0$ is again chosen so that $\textsf{Fairness}(\gamma)$ is between 0 and 1. Note that in the case of two groups, we have $z=\frac{1}{2}(x_{1}+x_{2})$ and
@@
with which is precisely the mutual fairness of Definition 3.1.
@@
We conclude with two remarks on this extension to $m$ groups. First, as in the case of two groups, there is no need to numerically solve optimal transport problems, as we provide a closed-form expression for the optimal transport problems. Second, we highlight that our extension to $m$ groups does not resort to the so-called multimarginal optimal transport problem, which might cause exponential complexity in the dimensionality.
@@ Appendix C Description and Properties of Datasets
To associate the notion of fairness developed in Sections 3.1 and 3.3 with the datasets and the outcomes from experiments in Section 3.2 and 4.2, we summarize the dataset statistics in Table 1. Minority Frac. is calculated as the fraction of the minority group nodes in the entire population. Fraction of Cross Edges evaluates heterophily in the dataset, by calculating the fraction of edges that connect different groups. A higher value means a more heterophilic network, whereas a lower value means a more homophilic network.
@@
A social network from Europe with $18,442$ nodes, where each node has a self-reported attributed gender (male or female). Men are the minority ($44.3\%$) and women are the majority ($55.6\%$). The data has moderate homophily.
@@
An interaction network from Instagram containing $553,628$ nodes, where everyone has a labeled gender ($45.57\%$ men and $54.43\%$ women). Each edge between two users represents a ‘like’ or ‘comment’ that one user gave another on a posted photo. The data has moderate homophily.

### v4

@@ 3.2 Mutual Fairness in Practice
We now investigate the use of our newly defined fairness metric across a variety of real-world datasets: Add Health (AH), Antelope Valley variants 00 to $23$ (AV_$\{0\text{-}23\}$[26], APS Physics (APS[14], Deezer (DZ[19], High School Gender (HS[15], Indian Villages (IV[1], and Instagram (INS[22]. Each dataset contains a social network with a chosen demographic partitioning the population into two groups (see Appendix C for details). We load the datasets as graphs $G(V,E)$. We then select a seedset $S$ of size $2$-$90$ (depending on the dataset) using the following heuristics: two group-agnostic seed selection strategies as our baselines, namely degree centrality (bas_d), and greedy (bas_g), proposed by Kempe et al. [12]. In addition, we implement two fair seed selection heuristics based on the equity metric, namely degree-central fair heuristic (hrt_d), and greedy fair heuristic (hrt_g), proposed by Stoica et al. [23]. To model the information spread, we use the Independent Cascade model (IC) for the diffusion of information [12] with a probability $p\in[0,1]$ for all edges. This process, being stochastic, is simulated $R$ times in a Monte Carlo sampling process to achieve $R$ final configurations (Definition 2.1) plotted together as a joint outreach distribution, in Fig. 3. Then we apply our distribution-aware notion of fairness from Section 3.1, mutual fairness. We keep $R=1,000$ throughout, but explore several values in $p,|S|$ (mentioned per experiment in the figures below) and exhaustively recorded with other hyperparameters in Appendix D. All details related to computational resources and development environment are available in Appendix G. The code for all our numerical experiments is available at https://github.com/nicolaslanzetti/fairness-sim-ot.
@@
As a first experiment, we study the joint outreach probability distribution for different datasets. We identify four qualitatively different outcomes, shown in Fig. 3 for a few of the datasets. Additional experiments with different propagation probability and seed selection strategies can be found in Appendix D. Fig. 3(a) is obtained on AH with bas_g selection strategy and $p=0.5,|S|=10$. We note how the joint outreach distribution is almost concentrated on the top right of the plane, i.e., the outcome is almost deterministic and highly fair and efficient. In turn, this trivializes both the expected value in the equity metric and the cost in the mutual fairness metric in Definition 3.1, which therefore essentially boils down to comparing the almost deterministic outreach fraction within each group. In these cases, our fairness metric does not provide additional insights. Such deterministic outcomes are typical of degree or greedy seedset outreach in dense graphs, such as AH, DZ, INS (refer to Appendix D), with extreme probability of conduction ($p\geq 0.5$ or $p\rightarrow 0$), and cross-group interconnectivity (see Table 1 in Appendix C). For moderate $p$ (e.g., $0.1$), the outreach probability distribution is concentrated along the diagonal (Fig. 3(b)). Thus, both the equity metric and our fairness measure are maximal. Nonetheless, our fairness metric provides additional insights: not only does the expected outreach within each group coincide, but also the outreach at every realization coincides (see the example in Section 3). Thus, our fairness metric provides a stronger certificate of fairness. As before, the same applies to AH, DZ, INS (see Appendix D). Intuitively, high cross-group interconnectivity in a dense graph already ensures fairness. Additionally, extreme $p$ values ensure deterministic outreach (either the information dies out at the seedset, or reaches everyone in the population). When propagation happens with moderate propagation probabilities, $p$, outreach appears as in Fig. 3(b). Fig. 3(c) represents APS for its hrt_g seedset outreach and $p=0.3,|S|=6$. Here, we observe a highly stochastic outcome, with many realizations for which almost no member of one group receives the information. Note that the phenomenon observed in Fig. 3(c) is the same as the one captured by our motivating example. We argue such an outcome should not be classified as fair, despite the expected value of the proportions being similar. Finally, Fig. 3(d) shows the AV_0 dataset with $p=0.3,|S|=4$, and bas_g selection strategy. We observe a more stochastic outreach compared to Fig. 3(b) with variance spread along, but not on the diagonal, with a small bias towards one group. Also in this case, both the equity and mutual fairness metrics characterize this outcome as fair, but mutual fairness is more informative as it requires outcomes to be fair at each realization.
@@
As a second experiment, we investigate the difference between mutual fairness and equity (difference in the expected value of the proportions), as a function of the conduction probability $p$. We consider the IV dataset as a case study and select seeds using bas_g. We show our results in Fig. 4. Our mutual fairness metric in Definition 3.1 shows a fundamentally different trend compared to the equity metric from Definition A.3. Importantly, for $p\in(0,0.5)$, both metrics have an opposite trend: equity fairness increases to some extent whereas our metric suggests a significant fall in fairness in this region. For $p\in(0.5,0.7)$, there is a decrease in equity fairness, while our fairness evaluation remains relatively constant. We notice similar trends for both metrics only for $p\in(0.8,1.0)$. The significant difference in the trend of the two metrics confirms our previous finding that mutual fairness is more informative than the equity metric and that the equity metric fails to adequately capture changes in fairness, see Sections 3.1 and 3.2. For more experiments on other datasets, we refer to Section D.2.
@@ 3.3 Trading off Fairness and Efficiency
In our setting, we can easily introduce the tradeoff in the transportation cost (2). Specifically, we can define the transportation cost as a weighted sum of the “diagonal distance” (measuring the difference in efficiency, dotted segment in Fig. 2) and the “orthogonal distance” (measuring the difference in fairness, solid segment in Fig. 2). Formally, for a given weight $\beta\geq 0$, the transportation cost reads
@@
The terms $1$ and $\sqrt{2}/\max\{1,2-2\beta\}$ in (4) ensure that the metric is non-negative and in $[0,1]$. Again, the optimal transport problem can be solved in closed form, which yields
@@ 4.1 Fairness-promoting Seed-selection Algorithm
Armed with a novel fairness metric, $\beta\mathrm{-}\textsc{Fairness}$, we now design an iterative seed-selection algorithm, which we call Stochastic Seedset Selection Descent (S3D), that strategically selects seeds taking into account all communities simultaneously. The pseudo-code is summarized in Algorithm 1. For its motivation and details, refer to Appendix E. For a given initial seedset, our algorithm explores new seeds and evaluates them on the efficiency-fairness metric $\beta\mathrm{-}\textsc{Fairness}$ as in (4) for a desired value of the fairness-efficiency tradeoff parameter $\beta$ (S3D_STEP() in Appendix E), to decide if the new seedset becomes a candidate for the optimized seedset. These seeds are searched for by iteratively sampling stochastically reachable nodes, up to a fixed depth, taken as a fraction of the graph diameter, from the current seedset (SEEDSET_REACH() in Appendix E) while making sure they contribute to a non-overlapping outreach (Algorithm 1::6-8). To avoid local minima of the generally non-convex objective, the procedure allows for visiting inferior seedsets on $\beta\mathrm{-}\textsc{Fairness}$ or even selecting completely random ones on rare occasions (Algorithm 1::12-18) using Metropolis Sampling [5]. Otherwise, a high $\beta\mathrm{-}\textsc{Fairness}$ encourages opting for the new seedset with high probability. Finally, we revisit all the seedset candidates collected so far and pick the one with the largest $\beta\mathrm{-}\textsc{Fairness}$ as the optimal seedset. For a sparse graph $G(V,E)$, with $E=O(V)$, choosing $|S|$ seeds, averaging over $R$ realizations to approximate outreach via Monte-Carlo sampling and exploring $k$ candidates using S3D_STEP suggests a total running time upper bound of $O(kR|S||V|)$ (see Appendix E for details). In practice, $k\in[500,1000],R=1000$ for $S\in[2,100]$ works well for all datasets.
@@ 4.2 Real-world Data
We test our algorithm across a variety of datasets (Appendix C) against our baselines (bas_d, bas_g). We initialize the S3D algorithm with the two baseline seedsets and hence include results from two separately optimized seedsets, S3D_d, S3D_g. Our results are shown in Fig. 6. Informally, we observe that our seed-selection mechanism “moves” the probability mass of the joint outreach probability distribution towards the diagonal, which ultimately increases the fairness of the resulting configuration. At the same time, efficiency either increases as well or suffers only a small decrease, as we investigate more in detail in our next experiment. Generally, datasets with high cross-group connections (AH, DZ, INS) yield moderately fair outreach with label-blind seed selection. Similarly, for datasets with low cross-group connections (APS) a label-blind strategy, in order to maximize efficiency, selects a diverse population of seeds from which all communities are reached. Therefore, label-blind algorithms work similarly to S3D. In other moderate cases (AV, HS, IV), instead, we observe significant improvements of S3D over label-blind strategies.
@@
In our final experiments, we compare several algorithms along with ours in terms of efficiency and mutual fairness across various datasets (see Appendix C). We consider the following algorithms: bas_d, bas_g, their fair heuristic counterparts, hrt_d, hrt_g, against our S3D_d, S3D_g, initialized via greedy and degree centrality baseline seeds, respectively. We show our results in Fig. 7. S3D achieves in almost all cases the highest fairness score ($y$-axis) and generally a slightly lower efficiency score ($x$-axis), compared to others. Thus, our seed-selection mechanism leads to fairer outcomes with only a minor decrease in efficiency.
@@
CEis high (>50%): The case where nodes interact more across groups than in their group was never observed. However, as long as the existence of edges does not reliably signal group information, we expect S3D to perform well based on a similar analysis.
@@ Appendix A Existing Fairness Metrics
$\bar{x}_{i}\coloneqq\frac{\mathbb{E}[|\{v\text{ reached }|v\in C_{i}\}|]}{|C_{i}|},\quad\forall i\in\{1,\dots,m\}.$
@@
$\frac{\mathbb{E}[|\{v\in S|v\in C_{i}\}|]}{|C_{i}|}=\frac{\mathbb{E}[|\{v\in S|v\in C_{j}\}|]}{|C_{j}|}\quad\forall i,j\in\{1,\dots,m\}.$
@@
Given a network with communities $C_{1},\dots,C_{m}$, a SIM algorithm that selects a seedset $S$ is said to be equitable if the algorithm propagation reaches all communities in a balanced way, i.e. $\bar{x}_{i}=\bar{x}_{j}$ for all $i,j\in\{1,\dots,m\}$.
@@ Appendix B Extension to Multiple Groups
We start with the definition of mutual fairness. We proceed as in Section 3.1 and define mutual fairness via optimal transport, which, in turn, requires defining a reference distribution and a transportation cost. The reference distribution is again the “ideal” distribution $\gamma^{\ast}=\delta_{(1,\ldots,1)}$ which encodes the case in which all members of all groups receive the information. As for the transportation cost, it suffices to generalize the transportation cost (2) to an $m$ dimensional space. Specifically, it can be defined as the distance between any given point $(x_{1},\ldots,x_{m})\in[0,1]^{m}$ in the hypercube and the diagonal line. For this, let
@@
$\textsc{Fairness}(\gamma)$ $=1-\alpha\mathbb{E}_{(x_{1},\ldots,x_{m})\sim\gamma}[\norm{z(x_{1},\ldots,x_{m})-(x_{1},\ldots,x_{m})}]$
@@
where the constant $\alpha>0$ is again chosen so that $\textsc{Fairness}(\gamma)$ is between 0 and 1. Note that in the case of two groups, we have $z=\frac{1}{2}(x_{1}+x_{2})$ and
@@
which is precisely the mutual fairness of Definition 3.1.
@@
We conclude with two remarks on this extension to $m$ groups. First, as in the case of two groups, there is no need to numerically solve optimal transport problems, as we provide a closed-form expression for the optimal transport problems. Second, we highlight that our extension to $m$ groups does not resort to the so-called multi-marginal optimal transport problem, which might cause exponential complexity in the dimensionality.
@@ Appendix C Description and Properties of Datasets
To associate the notion of fairness developed in Sections 3.1 and 3.3 with the datasets and the outcomes from experiments in Sections 3.2 and 4.2, we summarize the dataset statistics in Table 1. Minority % is calculated as the percentage of the minority group nodes in the entire population. Fraction of Cross Edges evaluates heterophily in the dataset, by calculating the fraction of edges that connect different groups. A higher value means a more heterophilic network, whereas a lower value means a more homophilic network.
@@
A social network from Europe with $18,442$ nodes, where each node has a self-reported attributed gender (male or female). Men are the minority ($44.3\%$) and women are the majority ($55.6\%$). The dataset has moderate homophily.
@@
An interaction network from Instagram containing $553,628$ nodes, where everyone has a labeled gender ($45.57\%$ men and $54.43\%$ women). Each edge between two users represents a ‘like’ or ‘comment’ that one user gave another on a posted photo. The dataset has moderate homophily.